# W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction

### v3

@@ Abstract
For a long time, in reconstructing 3D human bodies from monocular images, most methods opted to simplify the task by minimizing the influence of the camera. Using a coarse focal length setting results in the reconstructed bodies not aligning well with distorted images. Ignoring camera rotation leads to an unrealistic reconstructed body pose in world space. Consequently, the application scenarios of existing methods are confined to controlled environments. When confronted with complex and diverse in-the-wild images, they struggle to achieve accurate and reasonable reconstruction in world space. To address the above issues, we propose W-HMR, which decouples global body recovery into camera calibration, local body recovery, and global body orientation correction. We design the first weak-supervised camera calibration method for body distortion, eliminating dependence on focal length labels and achieving finer mesh-image alignment. We propose a novel orientation correction module to allow the reconstructed human body to remain normal in world space. Decoupling body orientation and body pose enables our model to consider the accuracy in camera coordinate and the reasonableness in world coordinate simultaneously, expanding the range of applications. As a result, W-HMR achieves high-quality reconstruction in dual coordinate systems, particularly in challenging scenes. Codes and demos have been released on the project page https://yw0208.github.io/w-hmr/.
keywords:
@@
Human mesh recovery, monocular 3D reconstruction, world space, camera calibration, motion capture
[…]
Qualitative results on SPEC-MTP [19]. The images in each triplet from left to right are the input image, the output in the camera coordinate, and the output in world coordinate. Compared to traditional methods, W-HMR shows obvious advantages for human mesh recovery in world space. W-HMR can efficiently correct erroneous poses in camera coordinate and ensure the reasonableness of poses in world space.
Introduction
Recovering 3D human meshes from monocular images is crucial in human-centered research, with wide-ranging applications in areas such as try-on, animation, and virtual reality. However, there is a noticeable lack of research on global body reconstruction from images captured under complex shooting conditions. For previous methods based on camera coordinate, the reconstructed bodies outside controlled environments are often unsuitable for real-world applications, as depicted in W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction.
@@ Introduction
Figure 1: Comparison among the camera models used by the other four models [14, 43, 19, 25] and our W-HMR. $f_{crop}$ and $f_{full}$ denote the focal lengths of the full and cropped images, respectively. $R_{c}$ is the camera rotation matrix. The red items mean trainable, i.e., predicted by neural networks.
@@
The primary challenge in this task arises from the limited camera information. Among camera parameters, focal length is critical in distortion effects, influencing the alignment of reconstructed 3D human bodies with original images. There is nothing inherently wrong with this misalignment. However, since most existing models are trained with 2D joint labels, this misalignment due to distortion affects the training and ultimately leads to decreased recovery accuracy. Additionally, camera rotation determines the reasonableness of the reconstructed body pose in the world (global) or camera (local) coordinate after coordinate transformation. Fig. 1 illustrates the treatment of the camera in existing models, emphasizing the need for improved methods. The focal length either involves simple manual definitions as seen in HMR [14] and CLIFF [25], or network predictions as utilized by SPEC [19] and Zolly [43]. As for camera rotation, most models [54, 53, 52, 20, 17, 18, 27] do not consider it and recover body in camera coordinate, except for SPEC. Even based on the most comprehensive camera model, SPEC still has limitations due to the inherent flaws in its CamCalib model, which is discussed as follows.
@@
Firstly, according to traditional methods like SPEC, the focal length is predicted based on background environmental information, which cannot handle finer distortions of body parts and is inadequate for the body recovery task. Moreover, focal length labels are always rare and not diverse enough, which can’t fit the need for model training. To address these challenges, we propose a weak-supervised camera calibration method. The core idea is that we don’t need to predict original focal lengths but the "reasonable" ones. Our method leverages body distortion information to predict a reasonable focal length for perspective projection, eliminating the need for supervision with rare focal length labels and making it more suitable for body tasks. Based on the full-perspective camera model, 2D supervision becomes more precise, thereby improving recovery accuracy.
@@
Secondly, the predicted camera rotation is not entirely accurate. With pseudo camera rotation as input, training a model like SPEC to predict poses in world coordinate would lead to error accumulation. The error in pseudo camera parameters would accumulate during training and impact the final recovery precision. Consequently, the accuracy of recovery is limited. Additionally, it decreases the model’s generalizability. For example, compared to SPEC’s advantages in world space, it performs poorly in camera coordinate, as shown in Table 2. We adopt an indirect approach for predicting body poses in world coordinate. First, we train our model to achieve optimal pose and shape accuracy in camera coordinate. Then, we propose a novel module named OrientCorrect to assist in correcting body orientation, thus avoiding the issues above and enabling human mesh recovery in world coordinate.
@@
    We present the first weak-supervised camera calibration method focus on the human recovery task. We achieve reasonable focal length prediction without focal length labels, but relying on human body distortion information. Better mesh-image alignment makes 2D supervision more precise and effectively improves recovery accuracy. Our method even outperforms the latest supervised method like Zolly [43].
@@
    We design a practical orientation correction module. The module solves the issue that previous methods cannot generate reasonable human poses in world space. This method is simple and effective and can be well transplanted to other models without destroying the original structure and performance.
@@
1 Related Work
1.1 Camera Model in Human Recovery
To visualize the reconstruction results and utilize 2D labels, 3D human recovery models require a camera model. The classical HMR [14] employed a weak-perspective camera model, which assumes an extremely large focal length of 5000 pixels to mitigate human distortion. This assumption is suitable for images captured with a long focal length when the human subject is distant from the camera. Due to its simplicity, the weak-perspective camera model has been widely adopted in subsequent works [54, 53, 18, 8, 17].
To the best of our knowledge, Kissos et al. [16] was the first to enhance the weak-perspective camera model in pose estimation. They assumed a constant camera Field of View (FoV) of 55 degrees instead of a fixed focal length. Later, CLIFF [25] also adopted the same camera model. SPEC [19] took a different approach by constructing a new dataset and training a camera parameter prediction model named CamCalib. The latest method, Zolly [43], indirectly predicts the focal length through a well-designed translation prediction module, albeit partially relying on focal length labels and only predicting the focal length of cropped images. Regarding camera rotation, it was usually overlooked in previous works. SPEC was the pioneering work to advocate for the recovery of the body in world coordinate for postural plausibility. It relied on CamCalib to predict camera parameters and used these parameters as model inputs to directly predict the global body. However, its one-step approach led to error accumulation, constraining reconstruction accuracy and application scope. In contrast, our decoupling of this process effectively addresses these issues.
1.2 Regression-based Method
In the 3D human body reconstruction domain, methods can be broadly classified into regression-based and optimization-based approaches. Optimization-based methods are beyond the scope of this paper and can be explored in the literature review [40]. Regression-based models can be further categorized into model-based and model-free approaches based on the output type. Model-based models refer to those that output template parameters, with the majority of current models [14, 8, 54, 17, 18] still relying on the SMPL algorithm [29], and a few [37, 58, 50, 51] based on SCAPE [2], GHUM&GHUML [47], and SMPL-X [36]. These models leveraged the prior knowledge embedded in parameter models but were limited to reconstructing finer mesh details due to constraints in degrees of freedom.
@@ 1.2 Regression-based Method
[…]
2 Method
The overview of W-HMR is shown in Fig. 2. The input image undergoes a PyMAF-like model to extract features and obtain the recovered human mesh in camera coordinates. We design a focal length predictor to predict the focal length implementing a full-perspective projection on the full image for 2D joint supervision. In addition, we propose the OrientCorrect module to correct the body orientation to keep the pose reasonable in the world space. Our method is introduced in five sections. We first present some prior knowledge in Section 2.1. This explains why we need to use the full perspective camera model and the full image information. In Section 2.2, we focus on how we predict the focal length and how we use the predicted focal length for supervision. In Section 2.3, we describe how OrientCorrect works, which is another important module. In Section 2.4, we describe how to stabilize model training without focal length labels. Finally, in Section 2.5, we add some introduction to more loss functions. To fully demonstrate the effectiveness of our method, we experiment with various network structures. Specific details can be found in Section 3.1.
2.1 Preliminary
@@
The whole process is how a 3D object is imaged on a 2D plane based on the perspective camera model. You may notice from Equ. 2 that when $f$ is set very large, the effect of $Z$ would become negligible and the perspective effect would disappear. The consequence is that the human mesh does not align well with the original image, making 2D supervision in error and ultimately affecting recovery accuracy.
@@
where $h_{bbox}$ refers to the height of the bounding box. For a more detailed reasoning process, please refer to [25] or [16]. Our goal is to leverage the relationship between $t_{b}^{z}$ and $f$ to indirectly predict $f$ by getting $t_{b}^{z}$, thus achieving full perspective.
2.2 Weak-supervised Camera Calibration
Focal length prediction is essentially an ill-posed problem. From Section 2.1, we can see too many influencing factors in the process from 3D objects to 2D images. Trying to predict the actual focal length through a neural network is almost impossible. So, we propose using human body distortion information to predict a "reasonable" focal length for full-perspective projection.
Camera Calibration The focal length predictor takes the final feature map $\varPhi_{3}$ as input and output a single value $t_{b}^{z}$. To constrain the range of the predicted $t_{b}^{z}$ [43], we employ a sigmoid function and scale it by 10. This adjustment is necessary because distortion becomes irrelevant when a person stands 10 meters from the camera. We add batch normalization to avoid gradient vanishing or exploding caused by the sigmoid function. The focal length can be calculated according to $f=\frac{sh_{bbox}t_{b}^{z}}{2}$, where the height of bounding box $h_{bbox}$ is known and scale $s$ is output by the PyMAF-like regression module. From Equ. 2, to implement the projection, we also need to know the translation of the body to the camera. $t_{b}^{z}$ has already been generated, so we need further to compute $(t_{b}^{x},t_{b}^{y})$. Following [16], $(t_{b}^{x},t_{b}^{y})=(t_{x},t_{y})$ when projecting in cropped image. But in the full image, the conversion functions for $(t_{x},t_{y})$ to $(t_{x}^{b},t_{y}^{b})$ are summarized below
@@ 2.2 Weak-supervised Camera Calibration
where $(c^{x}_{bbox},\,c^{y}_{bbox})$ is the center of the bounding box and $(W,H)$ is the width and height of the full image. The reason why projection is important in the human recovery task is that we need to use rich 2D joint labels to enhance supervision and improve model performance. By the way, we also explain why we keep emphasizing the full image here. For example, as shown in Fig. 3, the distortion of stretching to left and right in red circles can only be achieved by projecting on the full image rather than solely on cropped images. That’s why we emphasize the implementation of full perspective projection in the full image. Moreover, the actual body orientation cannot be correctly reflected in the cropped image when the human is not in the center of the original image. Therefore, full-image information, i.e., global information in our method, is also an important part of our model input.
@@
Weak Supervision Our main losses can be categorized into joint loss, vertex loss, and parameter loss. In this section, we only introduce the joint loss. Please refer to Section 2.5 for information on the others. The joint loss can be calculated using the following function:
@@
[…]
Figure 4: This figure depicts our model training paradigm, illustrating the detachment strategy at different stages to achieve stable model training. The symbol $\otimes$ represents detachment. Traditional 2D supervision refers to projecting 3D joints onto the cropped image to get 2D joints based on the weak-perspective camera model, just as traditional models did. The "freeze" means other modules are freezed in the third stage and only OrientCorrect are updated.
[…]
2.3 Orientation Correction
Our goal is to recover human meshes in world space. Weak-supervised camera calibration can only help the model to get the optimal mesh accuracy in camera coordinate. After that, we need to consider how to keep the human pose reasonable in world coordinate without destroying the original accuracy. In short, we need to predict the body orientation in world coordinate to replace the body orientation predicted in camera coordinate.
We propose the OrientCorrect for predicting body orientation in world coordinate. This module’s input comprises the camera rotation $R_{c}$ predicted by CamCalib, spatial body features $\phi_{3}^{s}$, global information, and the body orientation $\theta_{0}$ in camera coordinate. The $\theta_{0}$ is predicted by the PyMAF-like module. The output of OrientCorrect is the human body orientation in world coordinate. The specific procedure can be expressed as follows:

### v4

# Capturing Human Motion from Monocular Images in World Space with Weak-supervised Calibration

@@ Abstract
Previous methods for 3D human motion recovery from monocular images often fall short due to reliance on camera coordinates, leading to inaccuracies in real-world applications where complex shooting conditions are prevalent. The limited availability and diversity of focal length labels further exacerbate misalignment issues in reconstructed 3D human bodies. To address these challenges, we introduce W-HMR, a weak-supervised calibration method that predicts “reasonable” focal lengths based on body distortion information, eliminating the need for precise focal length labels. Our approach enhances 2D supervision precision and recovery accuracy. Additionally, we present the OrientCorrect module, which corrects body orientation for plausible reconstructions in world space, avoiding the error accumulation associated with inaccurate camera rotation predictions. Our contributions include a novel weak-supervised camera calibration technique, an effective orientation correction module, and a decoupling strategy that significantly improves the generalizability and accuracy of human motion recovery in both camera and world coordinates. The robustness of W-HMR is validated through extensive experiments on various datasets, showcasing its superiority over existing methods. Codes and demos have been made available on the project page https://yw0208.github.io/w-hmr/.
Index Terms:
@@
Human motion recovery, monocular 3D reconstruction, world space, camera calibration, motion capture
Figure 1: Given input images, we show recovered motion output by traditional methods based on the camera coordinate and our W-HMR based on the world coordinate. In contrast to traditional methods, W-HMR is capable of rectifying incorrect poses in the camera coordinate and guaranteeing the rationality of poses in the world space.
@@
Figure 2: Comparison among the camera models used by the other four models [1, 2, 3, 4] and our W-HMR. $f_{crop}$ and $f_{full}$ denote the focal lengths of the full and cropped images, respectively. $R_{c}$ is the camera rotation matrix, and $I$ is the identity matrix. $s$ means scale. $h_{bbox}$ and $h$ are the heights of the bounding box. But $h_{bbox}$ is dynamic, and $h$ is a constant value. $t_{b}^{z}$ refers to the vertical distance from the human body to the plane where the camera is located. The red items mean trainable, i.e., predicted by neural networks.
I Introduction
Recovering 3D human motion from monocular images is crucial in human-centered research, with wide-ranging applications in try-on, animation, and virtual reality. However, there is a noticeable lack of research on global body reconstruction from images captured under complex shooting conditions. For previous methods based on the camera coordinate, the reconstructed bodies outside controlled environments are often unsuitable for real-world applications, as depicted in Fig. 1.
@@ I Introduction
The primary challenge in this task arises from the limited camera information. Among camera parameters, focal length is critical in distortion effects, influencing the alignment of reconstructed 3D human bodies with original images. There is nothing inherently wrong with this misalignment. However, since most existing models are trained with 2D joint labels, this misalignment due to distortion affects the training and ultimately leads to decreased recovery accuracy. Additionally, camera rotation determines the reasonableness of the reconstructed body pose in the world (global) or camera (local) coordinate after coordinate transformation. Fig. 2 illustrates the treatment of the camera in existing models, emphasizing the need for improved methods. The focal length either involves simple manual definitions as seen in HMR [1] and CLIFF [4], or network predictions as utilized by SPEC [3] and Zolly [2].
As for camera rotation, most models [5, 6, 7, 8, 9, 10, 11, 12] do not consider it and recover body in the camera coordinate, except for SPEC. Even based on the most comprehensive camera model, SPEC still has limitations due to the inherent flaws in its CamCalib model, which are discussed below.
@@
Firstly, according to traditional methods like SPEC, the focal length is predicted based on background environmental information, which cannot handle finer distortions of body parts and is inadequate for the body recovery task. Moreover, focal length labels are always rare and not diverse enough, which can not fit the need for model training. To address these challenges, we propose a weak-supervised camera calibration method. The core idea is that we do not need to predict original focal lengths but the “reasonable” ones. Our method leverages body distortion information to predict a reasonable focal length for perspective projection. It eliminates the need for supervision with rare focal length labels and makes it more suitable for body tasks. Based on the full-perspective camera model, 2D supervision becomes more precise, improving recovery accuracy.
@@
Secondly, the predicted camera rotation is not entirely accurate. With pseudo camera rotation as input, training a model like SPEC to predict poses in the world coordinate would lead to error accumulation. The error in pseudo camera parameters would accumulate during training and impact the final recovery precision. Consequently, the accuracy of recovery is limited. Additionally, it decreases the model’s generalizability. For example, compared to SPEC’s advantages in world space, it performs poorly in the camera coordinate, as shown in Tab. II. We adopt an indirect approach for predicting body poses in the world coordinate. First, we train our model to achieve optimal pose and shape accuracy in the camera coordinate. Then, we propose a novel module named OrientCorrect to assist in correcting body orientation, thus avoiding the issues above and enabling human motion recovery in the world coordinate.
@@
    We present the first weak-supervised camera calibration method focus on the human recovery task. We achieve reasonable focal length prediction without focal length labels but rely on information about the human body distortion. Better mesh-image alignment makes 2D supervision more precise and effectively improves recovery accuracy. Our method even outperforms the latest supervised method like Zolly [2].
@@
    We design a practical orientation correction module. The module solves the issue that previous methods cannot generate reasonable human poses in world space. This method is simple and effective and can be transplanted to other models without destroying the original structure and performance.
@@
II Related Work
II-A Camera Model in Human Recovery
To visualize the reconstruction results and utilize 2D labels, 3D human recovery models require a camera model. The classical HMR [1] employed a weak-perspective camera model, which assumes a considerable focal length of 5000 pixels to mitigate human distortion. This assumption is suitable for images captured with a long focal length when the human subject is distant from the camera. Due to its simplicity, the weak-perspective camera model has been widely adopted in subsequent works [13, 7, 11, 14, 10].
To our knowledge, Kissos et al. [15] was the first to enhance the weak-perspective camera model in pose estimation. They assumed a constant camera Field of View (FoV) of 55 degrees instead of a fixed focal length. Later, CLIFF [4] also adopted the same camera model. SPEC [3] took a different approach by constructing a new dataset and training a camera parameter prediction model named CamCalib. The latest method, Zolly [2], indirectly predicts the focal length through a well-designed translation prediction module, albeit partially relying on focal length labels and only predicting the focal length of cropped images.
Regarding camera rotation, it was usually overlooked in previous works. SPEC was the pioneering work to advocate for the body’s recovery in the world coordinate for postural plausibility. It relied on CamCalib to predict camera parameters and used these parameters as model inputs to directly predict the global body. However, its one-step approach led to error accumulation, constraining reconstruction accuracy and application scope. In contrast, our decoupling of this process effectively addresses these issues.
II-B Regression-based Method
In the 3D human body reconstruction domain, methods can be broadly classified into regression-based and optimization-based approaches. Optimization-based methods are beyond the scope of this paper and can be explored in the literature review [16]. Regression-based models can be further categorized into model-based and model-free approaches based on the output type. Model-based models refer to those that output template parameters, with the majority of current models [1, 14, 13, 10, 11, 17] still relying on the SMPL algorithm [18], and a few [19, 20, 21, 22] based on SCAPE [23], GHUM&GHUML [24], and SMPL-X [25]. These models leveraged the prior knowledge embedded in parameter models but were limited to reconstructing finer motion details due to constraints in degrees of freedom.
@@ II-B Regression-based Method
Figure 3: Pipeline of W-HMR. A PyMAF-like backbone is adopted to extract features and predict the recovery results in the camera coordinate. ”glob info” refers to global information consisting of five elements. $(\varDelta c_{x},\,\varDelta c_{y})$ are offsets of the bounding box center relative to the center of the original image. FULL2 2D supervision refers to 2D joint supervision on full images, which is based on full-perspective camera model. The CamClib takes the whole image as input and outputs camera rotation matrices. The OrientCorrect takes three vectors as input and outputs body orientation in the world coordinate. $\{s,\,t_{x},\,t_{y}\}$ are scales and translations in the camera coordinate. $f$ is focal length and $(t_{b}^{x},\,t_{b}^{y},\,t_{b}^{z})$ is the translation to camera in the world coordinate. Note $\varPhi_{3}$ and $\phi_{3}^{s}$ are features extracted by the PyMAF-like backbone. If you are confused about our feature extraction, please refer to Sec. IV-A for our network details.
III Method
The overview of W-HMR is shown in Fig. 3. The input image undergoes a PyMAF-like model to extract features and obtain the recovered human motion in the camera coordinates. We design a focal length predictor to predict the focal length to implement a full-perspective projection on the full image for 2D joint supervision. In addition, we propose the OrientCorrect module to correct the body orientation to keep the pose reasonable in the world space. Our method is introduced in five sections. We first present some prior knowledge in Sec. III-A. This explains why we need to use the full perspective camera model and the full image information. In Sec. III-B, we focus on how we predict the focal length and how we use the predicted focal length for supervision. In Sec. III-C, we describe how OrientCorrect works, another essential module. In Sec. III-D, we describe how to stabilize model training without focal length labels. Finally, in Sec. III-E, we add some introduction to more loss functions. To fully demonstrate the effectiveness of our method, we experiment with various network structures. Specific details can be found in Sec. IV-A.
III-A Preliminary
@@
The whole process is how a 3D object is imaged on a 2D plane based on the perspective camera model. You may notice from Equ. 2 that when $f$ is set very large, the effect of $Z$ would become negligible, and the perspective effect would disappear. The consequence is that the human motion does not align well with the original image, making 2D supervision in error and ultimately affecting recovery accuracy.
@@
where $h_{bbox}$ refers to the height of the bounding box. This formula means that the 2-meter body range is imaged on the image, and the range becomes scale $s$ multiplied by the height of the bounding box $h_{bbox}$. Scale is why this formula is general. The scale is predicted by a neural network, which represents the size of the human in the bounding box. Therefore, it is dynamically adaptive, making this formula valid in most cases. The evidence is qualitative results in our paper. Our recovered mesh can be reprojected to overlap well with the human area. In addition, much work has adopted this assumption and this formula since it was proposed in 2018 [1]. For a more detailed reasoning process, please refer to [4] or [15]. Our goal is to leverage the relationship between $t_{b}^{z}$ and $f$ to indirectly predict $f$ by getting $t_{b}^{z}$, thus achieving full perspective.
III-B Weak-Supervised Camera Calibration
Focal length prediction is essentially an ill-posed problem. From Sec. III-A, we can see too many influencing factors in the process, from 3D objects to 2D images. Predicting the actual focal length through a neural network is almost impossible. So, we propose using human body distortion information to predict a ”reasonable” focal length for full-perspective projection.
Camera Calibration The focal length predictor takes the final feature map $\varPhi_{3}$ as input and output a single value $t_{b}^{z}$. Please refer to Sec.  IV-A for how to extract features $\varPhi_{3}$. To constrain the range of the predicted $t_{b}^{z}$ [2], we employ a sigmoid function and scale it by 10. This adjustment is necessary because distortion becomes irrelevant when a person stands 10 meters from the camera. We add batch normalization to avoid gradient vanishing or exploding caused by the sigmoid function. The focal length can be calculated according to $f=\frac{sh_{bbox}t_{b}^{z}}{2}$, where the height of bounding box $h_{bbox}$ is known and scale $s$ is output by the PyMAF-like regression module. From Equ. 2, to implement the projection, we also need to know the translation of the body to the camera. $t_{b}^{z}$ has already been generated, so we need further to compute $(t_{b}^{x},t_{b}^{y})$. Following [15], $(t_{b}^{x},t_{b}^{y})=(t_{x},t_{y})$ when projecting in cropped image. But in the full image, the conversion functions for $(t_{x},t_{y})$ to $(t_{x}^{b},t_{y}^{b})$ are summarized below
@@ III-B Weak-Supervised Camera Calibration
where $(c^{x}_{bbox},\,c^{y}_{bbox})$ is the center of the bounding box and $(W,H)$ is the width and height of the full image. Projection is important in human recovery because we need to use rich 2D joint labels to enhance supervision and improve model performance. By the way, we also explain why we keep emphasizing the full image here. For example, as shown in Fig. 4, the distortion of stretching to left and right in red circles can only be achieved by projecting on the full image rather than solely on cropped images. That’s why we emphasize the implementation of full perspective projection in the full image. Moreover, the actual body orientation cannot be correctly reflected in the cropped image when the human is not in the center of the original image. Therefore, full-image information, i.e., global information in our method, is also an important part of our model input.
@@
Weak Supervision Our main losses can be categorized into joint loss, vertex loss, and parameter loss. In this section, we only introduce the joint loss. Please refer to Sec. III-E for information on the others. The joint loss can be calculated using the following function:
@@
where $MSE(\cdot)$ represents mean square error. The $\land$ here denotes ground truth. The $J_{3D}$ denotes 3D joints, which are regressed from mesh vertices[1]. $J_{2D}$ signifies normalized 2D joint coordinate in the cropped image and ${J}_{2D}^{full}$ is normalized 2D joint coordinate in the full image. Note that 2D joints are the projected 3D joints on images. $J_{2D}$ and $J_{2D}^{full}$ are the same things but based on different coordinates (cropped image vs full image). The last term in the Equ. 5 is the key of our FULL2 2D supervision. It is worth mentioning that $\left(\frac{W}{w_{bbox}},\frac{H}{h_{bbox}}\right)$ is used to dynamically balance the weights and prevent the FULL2 2D loss from being too small, which could lead to ineffective supervision.
Figure 5: This figure depicts our model training paradigm, illustrating the detachment strategy at different stages to achieve stable model training. The symbol $\otimes$ represents detachment. Traditional 2D supervision refers to projecting 3D joints onto the cropped image to get 2D joints based on the weak-perspective camera model, just as traditional models did. The ”freeze” means other modules are frozen in the third stage, and only OrientCorrect is updated.
Weak supervision is reflected in this FULL2 2D supervision. We do not use any real focal length label to supervise the training of the focal length predictor. From Sec. III-A, we know that the focal length is involved in the 3D to 2D projection. So, by having 2D and 3D joint labels, we can let the predictor learn how to generate a correct focal length for 3D to 2D joint projection. After all, all this effort to make the projection more precise is so that the model can be trained more accurately. By the way, rich joint labels are easier to collect for model training than rare focal length labels.
III-C Orientation Correction
Our goal is to recover human motion in world space. Weak-Supervised camera calibration can only help the model to get the optimal mesh accuracy in the camera coordinate. After that, we must consider keeping the human pose reasonable in the world coordinate without destroying the original accuracy. In short, we need to predict the body orientation in the world coordinate to replace the body orientation predicted in the camera coordinate.
We propose the OrientCorrect for predicting body orientation in the world coordinate. This module’s input comprises the camera rotation $R_{c}$ predicted by CamCalib, spatial body features $\phi_{3}^{s}$, global information, and the body orientation $\theta_{0}$ in the camera coordinate. The $\theta_{0}$ is predicted by the PyMAF-like module. The output of OrientCorrect is the human body orientation in the world coordinate. The specific procedure can be expressed as follows: